The length of a rectangle is 6m longer than its width.
If the perimeter of the rectangle is 44m , find its area.

Answer by mananth(16946)

The length of a rectangle is 6m longer than its width.
If the perimeter of the rectangle is 44m , find its area.
..
width = x
length = x+6
..
2*(L+W)= perimeter
2*(x+x+6)=44
2*(2x+6)=44
/2
2x+6 = 22
2x=22-6
2x=16
/2
x= 8
width = 8 length = 14
Area = L*W
A= 14*8
A= 112 sq. m
...
m.ananth@hotmail.ca